# A Common Origin of Asymmetric Self-interacting Dark Matter and Dirac Leptogenesis

### v1

@@ IV Baryogenesis via Dirac Leptogenesis from $\eta$ decay
$\begin{split}\frac{dY_{\xi_{1}}}{dx}&=-\frac{x}{H(M_{\xi_{1}})}s<\sigma|v|_{(\xi_{1}\xi_{1}\rightarrow All)}>\left[Y_{\xi_{1}}^{2}-{Y_{\xi_{1}}^{eq}}^{2}\right]\\ &-\frac{x}{H(M_{\xi_{1}})}\Gamma_{(\xi_{1}\rightarrow All)}\left[Y_{\xi_{1}}-Y_{\xi_{1}}^{eq}\right]\\ \frac{dY_{L}}{dx}&=\epsilon_{L}\frac{x}{H(M_{\xi_{1}})}\Gamma_{(\xi_{1}\rightarrow All)}B_{L}\left[Y_{\xi_{1}}-Y_{\xi_{1}}^{eq}\right]\,,\end{split}$ (12)
@@
Fig. 4 depicts lepton asymmetry $Y_{L}$ and comoving number density of $\xi_{1}$, i.e., $Y_{\xi_{1}}$, as a function of the dimensionless parameter $x$. Coupling constants $f$ is taken as $10^{-4}$ and $\lambda$ is taken as $10^{-7}$. The typical value of the cross-section is taken as $\sigma|v|_{(\xi_{1}\xi_{1}\rightarrow All)}=10^{-25}\,{\rm GeV}^{-2}$. The solid red and blue curves show the evolution of number densities of $\xi_{1}$ and lepton asymmetry respectively. As temperature falls below the mass of $\xi_{1}$ (i.e., $x>1$), it decouples from the thermal bath (see Fig. 4). As $\xi_{1}$ starts decaying to $\ell^{-}\nu_{R}$, lepton asymmetry develops proportional to $B_{L}$ and $\epsilon_{L}$. Once the decay of $\xi_{1}$ is complete, the lepton asymmetry settles to a constant value. In Fig. 4, we have taken the branching ratio $B_{L}=0.5$, $\epsilon_{L}$ = $5\times 10^{-7}$ and the heavy scalar mass to be $M_{\xi_{1}}$ = $10^{10}$ GeV. Electroweak Sphalerons that violate $B+L$ (but conserve $B-L$) transfer a portion of the lepton asymmetry to a net baryon asymmetry $Y_{B}=-0.55Y_{L}$.
@@ V Asymmetric Dark matter from $\eta$ decay
CP-violating out of equilibrium decays $\xi_{1}^{-}\to\psi^{-}\chi,\,\xi_{1}^{+}\to(\psi^{-})^{c}\chi^{c}$ produce an asymmetry between $\chi$ and $\bar{\chi}$ as well as between $\psi$ and $\bar{\psi}$, similar to lepton asymmetry. The symmetric component of $\psi$ gets diluted by three annihilation processes: $\overline{\psi}\psi\to{\rm SM~{}SM},\,\overline{\psi}\psi\to\overline{\chi}\chi,\,\overline{\psi}\psi\to Z^{\prime}Z^{\prime}$. Note that, in presence of only the process $\overline{\psi}\psi\to{\rm SM~{}SM}$ i.e., via gauge couplings of $\psi$ to SM, $\psi$ remains under-abundant upto TeV scale, see Appendix A. If we assume that $\psi\psi Z^{\prime}$ coupling is of the same order as that of $\chi\chi Z^{\prime}$, the other two channels dominates over $\overline{\psi}\psi\to{\rm SM~{}SM}$, and the symmetric doublet component remains many orders below the correct relic ballpark for GeV scale DM. The asymmetric number density of $\psi$ gets converted to a net $\chi$ density through the decay process: $\psi\to\chi\bar{f}f$ (we assume $M_{\psi}>M_{\chi}$), induced via the soft $\mathcal{Z}_{2}$ symmetry breaking term $\mu^{2}H^{\dagger}\eta$. The symmetric part of DM $\chi$ annihilates into pairs of $Z^{\prime}$ as discussed above. Therefore, we can safely assume $\chi$ relic to be asymmetric, solely produced from CP-violating out of equilibrium decays $\xi_{1}^{-}\to\psi^{-}\chi,\,\xi_{1}^{+}\to(\psi^{-})^{c}\chi^{c}$. Feynman diagrams responsible for generation of asymmetries in dark sector are shown in Fig. 8.
@@ VI Dark Matter Self-interaction
[…]
@@
Cross-sections for each regime are listed in Appendix B. Perturbative calculations hold good only in the Born regime. In the resonant regime, for attractive potential we see quantum mechanical resonances for smaller velocities due to (quasi)-bound state formation Dutta:2022knf . Since in the present case, we have asymmetric DM with an repulsive potential, we do not see any resonance spikes. In Fig. 11, we show the self-interaction allowed parameter space for $g^{\prime}=0.1$ in $M_{\chi}$ $M_{Z^{\prime}}$ plane. We constrain $\sigma/M_{\chi}$ as:
@@
Figure 11: Parameter space for self-interaction in $M_{\chi}$ $M_{Z^{\prime}}$ plane with $g^{\prime}=0.1$.
@@ VII DM Direct Detection
We can constrain model parameters with data from direct search experiments like CRESST-III Abdelhameed:2019hmk and XENON1T Aprile:2018dbl . In Fig. 14, the most stringent constraints from CRESST-III (dotted contours in black) and XENON1T (dotted contours in pink) experiments are shown assuming $g^{\prime}=0.1$. For below (above) $10{\rm GeV}$ DM, we use constraints from CRESST-III (XENON1T) experiment. The region towards left of each contour is excluded for the mixing parameter mentioned alongside the curves.
VIII Conclusions
Neutrino oscillation experiments convincingly establishes non-zero neutrino mass. However, no consensus has been established regarding Dirac or Majorana nature for neutrinos, as experiments looking for neutrinoless double beta decay does not provide any conclusive evidence. In this work, we assume neutrinos to be Dirac particles. Then we explain baryon and dark matter asymmetry in a common framework.
@@ VIII Conclusions
We extend SM with a $SU(2)_{L}$ doublet $\psi$ and a singlet $\chi$. The singlet fermion $\chi$ being odd under the discrete $\mathcal{Z}_{2}$ symmetry behave as a candidate of DM. The same $\mathcal{Z}_{2}$ symmetry disallowed neutrino Dirac mass. $\mathcal{Z}_{2}$ symmetry is softly broken without destabilizing DM $\chi$, ensured by $M_{\chi}<M_{\psi}$. This also leads to a Dirac mass of neutrinos. A heavy scalar doublets ($\eta$), having non-trivial $\mathcal{Z}_{2}$ charge, undergoes out-of-equilibrium decay. Decay channels $\eta\to\nu_{R}\ell$ and $\eta\to\chi\psi$ generates lepton and dark matter asymmetries. As $B-L$ becomes an exact symmetry, the CP-violating decay of $\eta$ to $\nu_{R}\ell$ produces equal and opposite $B-L$ asymmetries in the left and right-handed sectors. The right-handed sector is weakly coupled to SM sector, required by neutrino Dirac mass. $B-L$ asymmetry in the left-handed sector converts to a net $B$-asymmetry via the $B+L$ violating sphaleron transitions. The asymmetries in the right-handed sector remains intact. The asymmetries neutralizes much after the electroweak phase transition when the sphaleron transitions gets suppressed. Meanwhile the symmetric DM component annihilates away to pairs of light mediators. The light mediators facilitates self-interaction among the DM particles, a essential feature for alleviating the small-scale $\Lambda{\rm CDM}$ anomalies. The kinetic mixing between dark and visible sectors facilitates direct detection of DM. We have found the relevant parameter space for SIDM confronted with direct search constraints from experiments like CRESST-III and XENON-1T. Since kinetic mixing suppresses all annihilations to SM final states, all gamma, cosmic, and neutrino fluxes are well below the near future reach of indirect probes Arina:2018zcq ; https://doi.org/10.1002/andp.19163561802 ; Cassel:2009wt ; Iengo:2009ni ; Slatyer:2009vg . The indirect signals of SIDM remains much below the current constraints by the experiments like Fermi-LAT Fermi-LAT:2013sme ; Fermi-LAT:2015att , MAGIC MAGIC:2016xys , HESS HESS:2018cbt , AMS-02 PhysRevLett.110.141102 , CMB by Planck Aghanim:2018eyx , and $\gamma$-rays by INTEGRAL Knodlseder:2007kh . Although $\chi$ is invisible at the collider, the doublet component $\psi^{\pm}$ can be pair produced in the collider  Bhattacharya:2015qpa ; Bhattacharya:2017sml , and their decay may lead to interesting signatures.
@@ Appendix A Relic Abundance of Doublet DM having only gauge interactions with the SM
If the DM is a vector-like fermion doublet $\psi$ of the form $\psi^{T}=(\psi^{0},\psi^{-})$ (with hypercharge $Y=-1$, where we use $Q=T_{3}+Y/2$), odd under an imposed $\mathcal{Z}_{2}$ symmetry, its neutral component becomes stable DM. The Lagrangian for the doublet DM is given by,
@@
where $\tau_{a}$ are Pauli spin matrices (generators of $SU(2)$), $g$ and $g_{Y}$ denote $SU(2)$ and $U(1)$ coupling strengths respectively. Expanding the covariant derivative of the Lagrangian given in Eq.20, we get the interaction term of $\psi^{0}$ and $\psi^{\pm}$ with the SM gauge bosons as follows:
@@
where $g=\frac{e}{\sin\theta_{W}}$ and $g_{Y}=\frac{e}{\cos\theta_{W}}$, $e$ and $\theta_{W}$ are the elementary charge and Weinberg angle respectively. All interactions being fixed by the SM gauge charge, the model does not have any free parameter except the mass $M_{\psi}$. So both relic density and direct search cross-sections become a function of DM mass alone. The model has been implemented in LanHEP SEMENOV1998124 and MicrOmegas Belanger:2013oya . The relic density obtained from the model as a function of DM mass is shown in Fig. 15. We observe that correct relic density is obtained only for DM mass around 1 TeV.

### v2

@@ IV Baryogenesis via Dirac Leptogenesis from $\eta$ decay
[…]
@@
Fig. 4 depicts lepton asymmetry $Y_{L}$ and comoving number density of $\xi_{1}$, i.e., $Y_{\xi_{1}}$, as a function of the dimensionless parameter $x$. Coupling constants $f$ is taken as $10^{-4}$ and $\lambda$ is taken as $10^{-7}$. The typical value of the cross-section is taken as $\sigma|v|_{(\xi_{1}\xi_{1}\rightarrow All)}=10^{-25}\,{\rm GeV}^{-2}$. The solid red and blue curves show the evolution of number densities of $\xi_{1}$ and lepton asymmetry respectively. As temperature falls below the mass of $\xi_{1}$ (i.e., $x>1$), it decouples from the thermal bath (see Fig. 4). As $\xi_{1}$ starts decaying to $\ell^{-}\nu_{R}$, lepton asymmetry develops proportional to $B_{L}$ and $\epsilon_{L}$. Once the decay of $\xi_{1}$ is complete, the lepton asymmetry settles to a constant value. In Fig. 4, we have taken the branching ratio $B_{L}=0.5$, $\epsilon_{L}$ = $5\times 10^{-7}$ and the heavy scalar mass to be $M_{\xi_{1}}$ = $10^{10}$ GeV. Electroweak Sphalerons that violate $B+L$ (but conserve BLB-L) transfer a portion of the lepton asymmetry to a net baryon asymmetry $Y_{B}=-0.55Y_{L}$.
Possible washout of the visible lepton asymmetry would require (i) chemical equilibration between the dark and visible sectors and (ii) an active lepton-number-violating (LNV) process. In our model, neither condition is realised. The only portal after $\eta$ decouples is the tiny kinetic mixing $\epsilon\lesssim 10^{-8}$, which ensures that, scattering rate of all relevant interactions is too slow compared to the Hubble rate. Besides, once the temperature drops below $T\ll M_{\xi_{i}}$, there is no operational LNV channel for any dark-visible process to couple to. Hence neither the dark nor the visible asymmetry is exposed to washout. A detailed wash-out analysis is provided in Appendix C.
@@ V Asymmetric Dark matter from $\eta$ decay
CP-violating out of equilibrium decays $\xi_{1}^{-}\to\psi^{-}\chi,\,\xi_{1}^{+}\to(\psi^{-})^{c}\chi^{c}$ produce an asymmetry between $\chi$ and $\bar{\chi}$ as well as between $\psi$ and $\bar{\psi}$, similar to lepton asymmetry. The symmetric component of $\psi$ gets diluted by three annihilation processes: $\overline{\psi}\psi\to{\rm SM~SM},\,\overline{\psi}\psi\to\overline{\chi}\chi,\,\overline{\psi}\psi\to Z^{\prime}Z^{\prime}$. Note that, in presence of only the process $\overline{\psi}\psi\to{\rm SM~SM}$ i.e., via gauge couplings of $\psi$ to SM, $\psi$ remains under-abundant upto TeV scale, see Appendix A. If we assume that $\psi\psi Z^{\prime}$ coupling is of the same order as that of $\chi\chi Z^{\prime}$, the other two channels dominates over $\overline{\psi}\psi\to{\rm SM~SM}$, and the symmetric doublet component remains many orders below the correct relic ballpark for GeV scale DM. The symmetric part of DM $\chi$ annihilates into pairs of $Z^{\prime}$ as discussed above. Therefore, we can safely assume $\chi$ relic to be asymmetric, solely produced from CP-violating out of equilibrium decays $\xi_{1}^{-}\to\psi^{-}\chi,\,\xi_{1}^{+}\to(\psi^{-})^{c}\chi^{c}$. Feynman diagrams responsible for generation of asymmetries in dark sector are shown in Fig. 8. The asymmetric number density of $\psi$ gets converted to a net $\chi$ density through the decay process: $\psi\to\chi\bar{f}f$ (we assume $M_{\psi}>M_{\chi}$), induced via the soft $\mathcal{Z}_{2}$ symmetry breaking term $\mu^{2}H^{\dagger}\eta$. This decay does not wash out the dark asymmetry, as it preserves the dark charge $D(\psi)=D(\chi)$ and therefore acts only as a transfer process that converts the primordial $\psi$ asymmetry into the stable $\chi$. The total dark asymmetry $Y_{\Delta D}=Y_{\Delta\psi}+Y_{\Delta\chi}$ is thus conserved exactly. We provide a detailed wash-out analysis considering all relevant interactions in Appendix C.
@@ VI Dark Matter Self-interaction
$\sigma_{T}=\int d\Omega(1-\cos\theta)\frac{d\sigma}{d\Omega}$ (18)
@@
Cross-sections for each regime are listed in Appendix B. Perturbative calculations hold good only in the Born regime. In the resonant regime, for attractive potential we see quantum mechanical resonances for smaller velocities due to (quasi)-bound state formation Dutta:2022knf . Since in the present case, we have asymmetric DM with an repulsive potential, we do not see any resonance spikes. In Fig. 11, we show the self-interaction allowed parameter space for $g^{\prime}=0.1$ in $M_{\chi}$ - $M_{Z^{\prime}}$ plane. We constrain $\sigma/M_{\chi}$ as:
@@
Figure 11: Parameter space for self-interaction in $M_{\chi}$ - $M_{Z^{\prime}}$ plane with $g^{\prime}=0.1$.
@@ VII DM Direct Detection
We constrain model parameters with data from direct search experiments like CRESST-III Abdelhameed:2019hmk and XENON1T Aprile:2018dbl . In Fig. 14, the most stringent constraints from CRESST-III (dotted contours in black) and XENON1T (dotted contours in pink) experiments are shown assuming $g^{\prime}=0.1$. For below (above) $10{\rm GeV}$ DM, we use constraints from CRESST-III (XENON1T) experiment. The region towards left of each contour is excluded for the mixing parameter mentioned alongside the curves.
VIII Collider Phenomenology
In this section we summarise the collider phenomenology of the dark fermion $\Psi$ and the dark gauge boson $Z^{\prime}$.
Absence of displaced vertices:
In the low-energy effective theory, the three-body decay $\psi\rightarrow\chi f\bar{f}$ is induced by integrating out the heavy scalar doublet $\xi_{i}$, generating the dimension-6 operator
$\mathcal{L}_{\rm eff}\supset\frac{y_{1}y_{2}}{M_{\xi_{i}}^{2}}(\bar{\chi}\,\Gamma\,\psi)(\bar{f}\,\Gamma\,f),$ (20)
where $f$ is a Standard Model fermion. Defining
$y_{\rm eff}^{2}\equiv y_{1}y_{2},\qquad\Lambda\equiv M_{\xi_{i}},$ (21)
the decay width is
$\Gamma_{\psi}\simeq\frac{y_{\rm eff}^{4}}{192\pi^{3}}\frac{M_{\psi}^{5}}{\Lambda^{4}},$ (22)
The proper decay length ($L_{0}=c\tau_{\psi}=\hbar c/\Gamma_{\psi}$) is obtained as,
$L_{0}\simeq 1.17\times 10^{13}\ \text{m}\;\left(\frac{1}{y_{\rm eff}}\right)^{4}\left(\frac{\Lambda}{10^{10}\ \text{GeV}}\right)^{4}\left(\frac{1\ \text{TeV}}{M_{\psi}}\right)^{5},$
so that even for $y_{\rm eff}=1$, $M_{\psi}=1~\mathrm{TeV}$ and $\Lambda=10^{10}~\mathrm{GeV}$ we find $L_{0}\simeq 1.17\times 10^{13}~\mathrm{m}$, far exceeding the size of any collider detector. For smaller $y_{\rm eff}$ or smaller $\psi$ mass, the lifetime (and hence the decay length) increases further. Thus throughout the entire region relevant for successful visible and dark asymmetry generation, $\psi$ is effectively stable at colliders; no displaced vertices are expected. The collider signature of $\psi$ is, therefore, that of a heavy long-lived particle. The charged component $\psi^{-}$ behaves as a heavy stable charged particle (HSCP), producing a slow, highly ionising track through the tracker and muon system, while the neutral component $\psi^{0}$ behaves as missing transverse energy (MET). The production channels, $pp\to\psi^{+}\psi^{-},pp\to\psi^{\pm}\psi^{0},pp\to\psi^{0}\psi^{0}$ thus lead to a characteristic mixture of HSCP tracks and MET, a hallmark signature of collider-stable electroweak doublets. Current CMS and ATLAS HSCP searches CMS:2010HSCP ; ATLAS:2019HSCP exclude long-lived charged fermions up to $1.3\text{--}1.6~\mathrm{TeV}$. Neutral components contribute to MET+ISR signatures, constrained by monojet searches ATLAS:2021monojet ; CMS:2021monojet , which give a weaker bound $M_{\psi}\gtrsim 100\text{--}200~\mathrm{GeV}.$
The dark gauge boson $Z^{\prime}$ is light ($M_{Z^{\prime}}\sim\mathrm{MeV}$) and the dark coupling is sizeable ($g^{\prime}\sim 0.1$), as required for self-interacting dark matter. However, the interactions of $Z^{\prime}$ with the Standard Model are controlled by the kinetic mixing, which is extremely small in the phenomenologically relevant region (see Fig. 14):
$\epsilon\lesssim 10^{-8},\qquad g^{\prime}_{\rm eff}=\epsilon\,g^{\prime}\sim 10^{-9}$
Such a light and ultra–feebly coupled vector is far below the sensitivity of all current LHC dilepton, monophoton, dijet, or invisible-$Z$ searches ATLAS:2019llZ ; CMS:2021monojet ; CMS:2021monophoton . Furthermore, as demonstrated in Appendix C, all portal-induced scatterings between dark and visible fermions satisfy
$\frac{\Gamma_{\rm scatt}}{H}\sim 10^{-20}\qquad\text{at }T=M_{\Psi},$ (23)
ensuring that the two sectors never thermalise after the epoch of asymmetry generation. This not only guarantees the preservation of the asymmetry, but also explains why the $Z^{\prime}$ does not produce observable collider signatures. Hence, in this set-up, displaced vertices and successful asymmetry generation are mutually incompatible, which is a distinctive feature of the model.
IX Conclusions
@@
As no consensus has been established regarding Dirac or Majorana nature for neutrinos yet, we, in this work, explain baryon and dark matter asymmetry in a common framework assuming assume neutrinos to be Dirac particles. We extend SM with a $SU(2)_{L}$ doublet $\psi$ and a singlet $\chi$. The singlet fermion $\chi$ being odd under the discrete $\mathcal{Z}_{2}$ symmetry behave as a candidate of DM. The same $\mathcal{Z}_{2}$ symmetry disallowed neutrino Dirac mass. $\mathcal{Z}_{2}$ symmetry is softly broken without destabilizing DM $\chi$, ensured by $M_{\chi}<M_{\psi}$. This also leads to a Dirac mass of neutrinos protected by a global BLB-L symmetry. A heavy scalar doublets ($\eta$), having non-trivial $\mathcal{Z}_{2}$ charge, undergoes out-of-equilibrium decay. Decay channels $\eta\to\nu_{R}\ell$ and $\eta\to\chi\psi$ generates lepton and dark matter asymmetries. As BLB-L is an exact symmetry, the CP-violating decay of $\eta$ to $\nu_{R}\ell$ produces equal and opposite BLB-L asymmetries in the left and right-handed sectors. The right-handed sector is weakly coupled to SM sector, required by neutrino Dirac mass. BLB-L asymmetry in the left-handed sector converts to a net $B$-asymmetry via the $B+L$ violating sphaleron transitions. The asymmetries in the right-handed sector remains intact. The asymmetries neutralizes much after the electroweak phase transition when the sphaleron transitions gets suppressed. Meanwhile the symmetric DM component annihilates away to pairs of light mediators. The light mediators facilitates self-interaction among the DM particles, a essential feature for alleviating the small-scale $\Lambda{\rm CDM}$ anomalies. The kinetic mixing between dark and visible sectors facilitates direct detection of DM. We have found the relevant parameter space for SIDM confronted with direct search constraints from experiments like CRESST-III and XENON-1T. Since kinetic mixing suppresses all annihilations to SM final states, all gamma, cosmic, and neutrino fluxes are well below the near future reach of indirect probes Arina:2018zcq ; https://doi.org/10.1002/andp.19163561802 ; Cassel:2009wt ; Iengo:2009ni ; Slatyer:2009vg . The indirect signals of SIDM remains much below the current constraints by the experiments like Fermi-LAT Fermi-LAT:2013sme ; Fermi-LAT:2015att , MAGIC MAGIC:2016xys , HESS HESS:2018cbt , AMS-02 PhysRevLett.110.141102 , CMB by Planck Aghanim:2018eyx , and $\gamma$-rays by INTEGRAL Knodlseder:2007kh .
Although direct production of either $\psi$ or $Z^{\prime}$ at high-energy colliders is highly suppressed, the model exhibits several distinctive phenomenological features. Both the visible and dark asymmetries originate from the same CP-violating decays of the heavy scalar doublets, implying correlated phases and a common dynamical origin of baryogenesis and dark-genesis. The dark sector contains a self-interacting MeV-scale vector mediator $Z^{\prime}$ with an ultra-feeble kinetic portal to the Standard Model, while the heavy dark fermion $\psi$ is effectively stable on detector scales. This characteristic combination of ‘self-interacting dark matter + ultra-feeble portal + collider-stable $\psi$ is not realised in typical ADM or leptogenesis scenarios and therefore provides a distinguishing phenomenological signature of the framework. At future high-energy colliders such as the ILC ILC:2013 , CLIC CLIC:2018 , FCC-hh FCC:2019 , CEPC CEPC:2018 , and SPPC SPPC:2015 , the pair-production of $\psi$ can be significantly enhanced. the charged component of $\psi$ would appear as a heavy stable charged particle (HSCP), whereas the neutral component would manifest as a large missing-energy signal. Although $Z^{\prime}$ remains inaccessible due to its tiny kinetic mixing, the observation of long-lived $\psi$ states, together with astrophysical measurements of dark self-interactions would provide complementary avenues to probe the underlying dark sector and the unified origin of visible and dark asymmetries in this model.
@@ Appendix A Relic Abundance of Doublet DM having only gauge interactions with the SM
If the DM is a vector-like fermion doublet $\psi$ of the form $\psi^{T}=(\psi^{0},\psi^{-})$ (with hypercharge Y=1Y=-1, where we use $Q=T_{3}+Y/2$), odd under an imposed $\mathcal{Z}_{2}$ symmetry, its neutral component becomes stable DM. The Lagrangian for the doublet DM is given by,
@@
where $\tau_{a}$ are Pauli spin matrices (generators of $SU(2)$), $g$ and $g_{Y}$ denote $SU(2)$ and $U(1)$ coupling strengths respectively.
Expanding the covariant derivative of the Lagrangian given in Eq.24, we get the interaction term of $\psi^{0}$ and $\psi^{\pm}$ with the SM gauge bosons as follows:
@@
where $g=\frac{e}{\sin\theta_{W}}$ and $g_{Y}=\frac{e}{\cos\theta_{W}}$, $e$ and $\theta_{W}$ are the elementary charge and Weinberg angle respectively.
All interactions being fixed by the SM gauge charge, the model does not have any free parameter except the mass $M_{\psi}$. So both relic density and direct search cross-sections become a function of DM mass alone. The model has been implemented in LanHEP SEMENOV1998124 and MicrOmegas Belanger:2013oya . The relic density obtained from the model as a function of DM mass is shown in Fig. 15. We observe that correct relic density is obtained only for DM mass around 1 TeV.